If i spent 8.39 in canadian dollars and my credit card charged me US$ 7.66 What is the the currency rate? Please expalin. Thanks MP
Answer by checkley71(8403)

8.39/7.66=X/1 OR 7.66X=8.39 OR X=1.0953 OR 1.0953 CANADIAN DOLLARS FOR EACH US DOLLAR.